OpenStudy (anonymous):

what is b^2-2b-2 simplified

OpenStudy (anonymous):

b(b-2)-2 or (b-1)^2 - 3

OpenStudy (unklerhaukus):

\[b^2-2b-2=0\] \[b_{1,2}=\frac{-(-2)\pm\sqrt{(-2)^2-4(1)(-2)}}{2(1)}\\=\frac{2\pm\sqrt{4+8}}{2}\\=\frac{2\pm\sqrt{12}}{2}\\=\frac{2\pm2\sqrt{3}}{2}\\=1\pm\sqrt{3}\] \[\Big(b-(1+\sqrt 3)\Big)\Big(b-(1-\sqrt 3)\Big)=0\] \[b^2-2b-2=\Big(b-(1+\sqrt 3)\Big)\Big(b-(1-\sqrt 3)\Big)\]